• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name bogdanw/V. Sulkosky

    Log entry time 08:05:12 on September11,2009

    Entry number 289081

    Followups:

    keyword=Owl shift summary

    Full beam ~ 100 muA at 00:50
    check spot++ done at 1:30
    01:30-04:05 Beam tuning and profile check.
    04:05-04:15 Raster check. New size 11 mm x 7.5 mm.
    04:16 Move Target to LH2 loop 3.
    04:22 Ask for 100 uA for production.
    04:39 Move target to optics while MCC continues tuning.
    06:10 Beam tuning complete. MCC now adjusting beam position to
    nominal production settings.
    06:32 Beam profile checked with harps, size is comparable to earlier
    results. Asked MCC to restore raster.
    06:40 Raster check completed 5 mm x 4.6 mm.
    06:40 Move target to loop 3 LH2.
    07:20 Raster size changed to 11 mm x 8.5 mm.
    Lumi asymmetries still ~ 1500 ppm.


    Run Start Events Comments
    ++++++++++++++++++++++++++++++++++++++++++++++++++++++++
    12224 06:50 55.3k Lumi RMS ~ 2000 ppm
    12225 07:21 Continue to next shift, Raster 11 mm x 8.5 mm